# Bayesian Functional Graphical Models with Change-Point Detection

### v1

@@ 5 Conclusions
Although the sea surface temperature data could be modeled accurately using a relatively small number of basis functions ($K=10$), other applications may require larger values of $K$. However, increasing $K$ significantly increases the computational burden, since the relevant precision matrices are $(pK)\times(pK)$. In general, the computational cost of our proposed model is dominated by the number of graphs and the graph size $(pK)\times(pK)$. Alternatively, neighborhood selection approaches may provide more scalability for these settings (Zhao et al.,, 2021), and similarly could be extended to incorporate (one or more) changepoints for dynamic graph estimation. The model performance benefits from estimating each edge in $\bm{G}$ separately using marginal edge inclusion probabilities, which is standard in Bayesian graphical models.
@@ Appendix A Details of the MCMC algorithm
Appendix B Supplementary material for prior evaluation
@@
Table 2: Prior edge inclusion probability of graph $G$ in the coefficient space using different values of $K$ and $h$. Parameters are set to $\alpha=2$, $\beta=7$, $v_{0}=2$ and $p=15$.
@@
Figure 6: Prior distributions of an element in the precision matrix given edge = 0 (left) and edge = 1 (right). The scales of the x-axes are different.
@@
Figure 7: Distribution of posterior block-wise Frobenious norm of the precision matrix given edge estimation in the functional space. The left plot is when the corresponding edge in the functional space is estimated to be 0; the right plot is when the corresponding edge in the functional space is estimated to be 1.

### v2

@@ 5 Conclusions
Although the sea surface temperature data could be modeled accurately using a relatively small number of basis functions ($K=10$), other applications may require larger values of $K$. However, increasing $K$ significantly increases the computational burden, since the relevant precision matrices are $(pK)\times(pK)$. In general, the computational cost of our proposed model is dominated by the number of graphs and the graph size $(pK)\times(pK)$. Alternatively, neighborhood selection approaches may provide more scalability for these settings (Zhao et al.,, 2024), and similarly could be extended to incorporate (one or more) changepoints for dynamic graph estimation. The model performance benefits from estimating each edge in $\bm{G}$ separately using marginal edge inclusion probabilities, which is standard in Bayesian graphical models.
@@ Appendix A Details of the MCMC algorithm
Appendix B Supplementary material for simulation study
@@
Table 2: Prior edge inclusion probability of graph $\bm{G}$ in the coefficient space using different values of $K$ and $h$. Parameters are set to $\alpha=2$, $\beta=7$, $v_{0}=2$ and $p=15$.
@@
Figure 6: Prior distributions of an element in the precision matrix given that the edge is absent (left) or present (right). The scales of the x-axes are different.
@@
Figure 7: Distribution of posterior block-wise Frobenious norm of the precision matrix given edge estimation in the functional space. The left plot is when the corresponding edge in the functional space is estimated to be absent; the right plot is when the corresponding edge in the functional space is estimated to be present.
Figure 8: Prior edge inclusion probability for $\mathcal{G}$ (y-axis) versus mean of the Beta prior, $\alpha/(\alpha+\beta)$, (x-axis) for $p=5$ (top left), $p=10$ (top right), $p=15$ (bottom left), and $p=20$ (bottom right). The edge inclusion probabilities are uniformly higher than for $\bm{G}$ (Figure 1).